Secret:

Give the values of a, b, and c needed to write the equation's standard form. 1/4x^2 + 5 = 0 A = 1/4; B = 5; C = 0 A = 1; B = 0; C = 20 A = 1; B = 0; C = -5

Hero:

@Secret, you are given \(\dfrac{1}{4}x^2 + 5x = 0\). One thing you should know is that an equation in standard form cannoat have any fractions as coefficients. So therefore what number can we multiply both sides of the equation by to get rid of the fraction?

Secret:

5x?

Hero:

@Secret, have you tried doing that? If so, let me know what you get afterwards.

Secret:

Is it 5x?

jhonyy9:

no not is 5x do you remember how you assume two fractions ? \[\frac{ 1 }{ 2 }x + \frac{ 2 }{ 3 }x = ?\] so in this your case there are \[\frac{ 1 }{ 4 }x^2 + 5x = 0 \] do you know how many is the denominator of 5x ?

Hero:

@Secret can you tell me what the result will be if you multiplied both sides of the equation by 4?